OpenStudy (anonymous):

(7, -2) (26, 15) write in point slope form

OpenStudy (anonymous):

you got the slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

when i tried to get the slope I got 17/19 and it said it was wrong

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

\[\frac{17}{19}\]is right for the slope

OpenStudy (anonymous):

maybe you just weren't done i think it is asking for the equation of the line

OpenStudy (anonymous):

i got y-15=17/19(x-(-2)) for point slope form and it said it was wrong

OpenStudy (anonymous):

ooh i see the problem

OpenStudy (anonymous):

you have to pick one point, not use the y coordinate of the different points

OpenStudy (anonymous):

pick this one \((7, -2)\) as \((x_1, y_1)\) or pick \((26, 15)\) either way, but pick one

OpenStudy (anonymous):

ohh okay i see. thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

how do I multiply 17/19 by 26